OpenStudy (anonymous):

If r>0 and r^t=6.25r^t+2, then r =?

OpenStudy (anonymous):

what have you tried?

OpenStudy (anonymous):

Nothing i dont even know how to start the problem

OpenStudy (anonymous):

move all the like terms to one side of the equation

OpenStudy (anonymous):

err... is this r^t = 6.25r^(t+2) or r^t = 2+6.25r^t

OpenStudy (anonymous):

is the exponent on the r on the right hand side 't' or 't+2' ?

OpenStudy (anonymous):

its the first way that you had it

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

yes its r^t+2

OpenStudy (anonymous):

ok this is pretty trivial then... use the rule \[x ^{a+b} = x ^{a}x ^{b}\]

OpenStudy (anonymous):

on \[r ^{t+2}\]

OpenStudy (anonymous):

What do i do after that

OpenStudy (anonymous):

\[r ^{t} =6.25r ^{2}r ^{t}\]

OpenStudy (anonymous):

r^t is a factor on both sides of the equation...

OpenStudy (anonymous):

\[1=6.25r ^{2}\]

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

ok but what happens to the r^2

OpenStudy (anonymous):

solve for r now:)

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes thank you!!

OpenStudy (anonymous):

sure!